OpenStudy (anonymous):

convert a line with slope=4 and point (-2,-13) to slope intercept form

OpenStudy (campbell_st):

slope intercept form is y = mx + b m = the slope and b is the y-intercept substitute it the slope m = 4 the use the point x = -2 and y = -13, substitute them and find the value of b then put that back into the equation.

OpenStudy (anonymous):

so it would be y=-8+-13?

OpenStudy (campbell_st):

not quite -13 = 4*(-2) + b

OpenStudy (anonymous):

okay so would we do -13=-8 and do add 13 to -8?

OpenStudy (campbell_st):

well its -13 = -8 + b whats the value of b

OpenStudy (anonymous):

-5?

OpenStudy (campbell_st):

so now you need the equation in the form y = mx + b you know m and just found b... substitute them for the answer

OpenStudy (anonymous):

4+-5=-1 so -1 is the answer?

OpenStudy (campbell_st):

nope leave x and y as letters just put in m = 4 and b = -5 what do you think it might be

OpenStudy (anonymous):

y=4x+-5?

OpenStudy (campbell_st):

thats it.... but you only need 1 operator y = 4x - 5

OpenStudy (anonymous):

thank you soo much!! :)